OpenStudy (anonymous):

how do you solve a rational equation symbolically? (x)/(x+12)-2=3

OpenStudy (anonymous):

turns out that solving it symbolically is solving it "normally" lol. first step is to add 2 to both sides: \[\frac{ x }{ x+12 } = 5\] then to multiply both sides by (x + 12) \[x = (x + 12)*5\] \[x = 5x + 60\] subtract x - 60 from both sides: \[4x = -60\] divide both sides by 4: x = -15

OpenStudy (anonymous):

why then say "graphically"-weird- i got lost with subtracting x-60.... how do you get to 4x=-60? :)

OpenStudy (anonymous):

I meant "symbolically" not graphically...ugh, late and doing homework-lol

OpenStudy (anonymous):

from: x = 5x + 60 i subtracted 60 from both sides x - 60 = 5x then i subtracted x from both sides -60 = 4x combined it to one step to save time :P

OpenStudy (anonymous):

oh. it's my mistake. technically its subtracting (x + 60) since subtracting (x - 60) gives 60 - x when the intention is - 60 - x. i'll stick to one step at a time :P

OpenStudy (anonymous):

ahh...i see , thank you so much!!

OpenStudy (anonymous):

glad i could help :)